### v1

@@ 7 Conclusion
This study introduced a QCAE model that leverages the power of QAOA to seamlessly substitute for the conventional AE, serving as the latent space. This substitution improves data learning in a higher-dimensional space. We demonstrated the efficiency of our proposed QCAE method through its application to MNIST image denoising. The proposed QCAE method exhibited superior denoising performance, as it achieved lower training loss and increased SSIM, compared to the CCAE method—a finding validated by experimental results—demonstrating its superior capacity to preserve image fidelity. The developed model maintained its superior denoising capabilities when applied to the MNIST dataset. Furthermore, we examined QAOA performance to determine its effectiveness across diverse circuit configurations and layer variations. Finally, we conducted an experiment using PSR optimization to evaluate its impact on QAOA circuit accuracy. This investigation underscores the crucial role of optimization in achieving desirable outcomes. However, we remark that state-of-the-art and out-of-the-box classical algorithms cannot be easily outperformed by current quantum computers and algorithms. The potential future research directions are as follows: First, we aim to evaluate the effectiveness of our proposed approach by conducting tests on real-world practical images and providing evidence to validate its viability. Second, we aim to investigate how different optimization methods affect the performance of our proposed model on the MNIST dataset. Furthermore, given that notice noise in real quantum machines can significantly impact the outcomes of the method, employing noise mitigation techniques is required to address and alleviate these effects.
@@
[…]

### v2

@@ 7 Conclusion
This study introduced a QCAE model that leverages the power of QAOA to seamlessly substitute for the conventional AE, serving as the latent space. This substitution improves data learning in a higher-dimensional space. We demonstrated the efficiency of our proposed QCAE method through its application to MNIST image denoising. The proposed QCAE method exhibited superior denoising performance, as it achieved lower training loss and increased SSIM, compared to the CCAE method—a finding validated by experimental results—demonstrating its superior capacity to preserve image fidelity. The developed model maintained its superior denoising capabilities when applied to the MNIST dataset. Furthermore, we examined QAOA performance to determine its effectiveness across diverse circuit configurations and layer variations. Finally, we conducted an experiment using PSR optimization to evaluate its impact on QAOA circuit accuracy. This investigation underscores the crucial role of optimization in achieving desirable outcomes. However, we remark that state-of-the-art and out-of-the-box classical algorithms cannot be easily outperformed by current quantum computers and algorithms. Notably, we did not include the mixer QAOA typically used in the QAOA circuits. The reason for this omission is that our primary focus was on evaluating the effectiveness of the QAOA mixing Hamiltonian in enhancing the latent space representation within the QCAE model. In standard QAOA, the mixing Hamiltonian is indeed a fundamental component and is typically chosen as a transverse field. This Hamiltonian helps to explore the solution space by inducing transitions between different quantum states. By isolating the mixing Hamiltonian, we aimed to specially assess its impact on denoising performance without the added complexity introduced by the mixer QAOA. The potential future research directions are as follows: First, we aim to evaluate the effectiveness of our proposed approach by conducting tests on real-world practical images and providing evidence to validate its viability. Second, we aim to investigate how different optimization methods affect the performance of our proposed model on the MNIST dataset. Furthermore, given that notice noise in real quantum machines can significantly impact the outcomes of the method, employing noise mitigation techniques is required to address and alleviate these effects. Future work will also explore how the dimensionality of the latent space influences the QCAE’s representational capacity and learning performance, which could provide insights for optimizing the balance between model complexity and denoising accuracy.
@@
CRediT authorship contribution statement
Tara Kit: Conceptualization, Methodology, Validation, Visualization, Software, Writing – review & editing. Kimsay Pov and Kimleang Kea: Writing - review & Validation. Won-Du Chang: Writing – review, Validation, Visualization. Hee Chul Park: Writing – review, Validation. Youngsun Han: Supervision, Validation, Visualization, Writing – review & editing.